earth's radius is approximately 4000 miles. about 2/3 of earth's surface is covered by water. estimate the land area on earth?
Answer by checkley77(12844)

A=4PIR^2
A=4*3.12*4000^2
A=199,680,000 SQ. MILES TOTAL AREA.
1/3A=199,680,000/3
A=66,493,440 SQ. MILES OF LAND,